# Lite Any Stereo: Efficient Zero-Shot Stereo Matching

### 16 Mar 2026 v2

@@ Abstract
Figure 1: Zero-shot prediction on in-the-wild stereo images. The proposed method achieves accurate disparity estimation across diverse scenarios and maintains high efficiency, even on older-generation GPUs.
@@ 4.4 Evaluation
Inference Time Analysis. Prior works report inference times measured on different hardware, making direct comparisons unreliable. To ensure fairness, we benchmark all methods under the same hardware setup. In Tab. 7, our method consistently achieves the fastest speed across different GPUs, highlighting its superior efficiency. Notably, it requires only 2.5 GB of GPU memory for 2K inputs, indicating strong potential for deployment on embedded devices, drones, and other resource-constrained platforms.
@@ 4.5 Limitations
Table 7: Inference time (ms) evaluated locally on different GPU hardware under the same experimental setup.
Method 1080 4090 A5000 A100
Lite-CREStereo++ [25] 63 38 55 40
LightStereo-M [20] 24 22 27 21
LightStereo-L [20] 34 30 35 27
BANet-2D [66] 58 30 63 27
BANet-3D [66] 40 25 44 22
Lite Any Stereo 21 19 23 17

### 15 Apr 2026 v3

@@ Abstract
Figure 1: Zero-shot predictions on in-the-wild stereo images. The proposed method produces accurate disparity estimates across diverse scenarios while maintaining high efficiency, requiring less than 1% of the computational cost than accuracy-focused methods.
@@ 4.4 Evaluation
Inference Time Analysis. Since prior works often report inference times on different hardware, direct comparisons can be unreliable. To ensure a fair comparison, we benchmark all methods locally under the same experimental setup across multiple GPU platforms. As shown in Tab. 7, our method ranks among the fastest across different GPUs, demonstrating strong computational efficiency. These results suggest its practical potential for deployment on embedded devices and other resource-constrained platforms.
@@ 4.5 Limitations
Table 7: Inference time (ms) measured locally on different GPUs under the same experimental setup (KITTI resolution), with torch.compile enabled and CUDA synchronization applied.
Method 4090 A5000 A100 H200
Lite-CREStereo++ [25] 14 30 20 16
LightStereo-M [20] 4 8 6 5
LightStereo-L [20] 7 16 10 7
BANet-2D [66] 4 8 7 5
BANet-3D [66] 6 12 10 7
Lite Any Stereo 4 9 7 6